On a sunny day, a rooftop solar panel delivers 75 W of power to the house at an emf of 17 V 

Part A 

How much current flows through the panel?
Answer #1

Part A: Electric power \(\mathrm{P}=\mathbb{I} \Delta V\)

Current through the panel \(\mathrm{I}=\mathrm{P} / \Delta V\)

\(\mathrm{I}=75 / 17=4.41\) amperes